# Beyond the Leaderboard: Rethinking Medical Benchmarks for Large Language Models

### v1

@@ 4.4 Guideline for Benchmark Validity & Performance Verification
Findings: The Risk of Misdirected Progress. Formal validation of a benchmark’s own scientific properties is rare. While most benchmarks can differentiate models, only 57% (30 of 53) provide a compelling analysis of their content validity, and just 40% (21 of 53) are grounded in scenarios with high real-world authenticity. Without proper validation, a benchmark’s results are difficult to interpret. Poor content validity creates blind spots, while poor construct validity means the benchmark may not measure the claimed capability at all, leading to misleading conclusions and misdirected research efforts (Alaa et al., 2025). This lack of self-reflection compromises the scientific value of the work built upon these tools, risking optimization for metrics that lack genuine clinical utility.
@@ 4.5 Guideline for Documentation, Openness, & Governance
Findings: A Fragmented and Unsustainable Ecosystem. While most benchmarks provide public access to their assets (52 of 53), the governance required for sustainable impact is critically lacking. A significant 42% (22 of 53) do not specify a usage license, creating barriers to adoption. Most critically, 85% (45 of 53) have no clear maintenance plan, and 66% (35 of 53) lack a public feedback channel. This effectively renders them "fire-and-forget" artifacts destined for obsolescence. This practice fosters a fragmented and unsustainable ecosystem, where effort is wasted on disposable artifacts rather than on building a lasting, reliable infrastructure for scientific progress (Arvan et al., 2022).
@@ 5 Discussion
Our analysis of 53 medical LLM benchmarks, viewed through the MedCheck framework, reveals a field at a critical crossroads. While research is active, systemic deficiencies threaten the validity and utility of this work.
@@ 5.1 Implications: The Urgent Need for a Paradigm Shift
Continuing with the status quo is not only scientifically unsound but also clinically irresponsible. The current trajectory of medical LLM benchmarking fosters an "illusion of progress," where leaderboard scores climb on contaminated or clinically irrelevant tasks, masking a lack of genuine advancement in capabilities required for real-world application (Laskar et al., 2024; Chang et al., 2023). This misdirects research efforts, misinforms stakeholders, and delays the responsible integration of AI into healthcare by creating brittle, biased systems that could harm patients (Obermeyer et al., 2019).
@@ 5.2 Our Guideline as a Catalyst for Change
The five-phase lifecycle and 46-criteria checklist presented in this paper provide a practical toolkit for this essential shift. Our framework offers a clear, actionable roadmap for developers to create higher-quality benchmarks and for users to critically assess the limitations of existing ones. Adopting this framework can directly mitigate the critical issues identified in our study by promoting:
Clinical Grounding: Mandating expert involvement and alignment with medical standards from Phase I.
Data Integrity: Enforcing transparent sourcing, quality control, and contamination checks in Phase II.
Safety-Oriented Evaluation: Requiring the assessment of reasoning, robustness, and uncertainty in Phase III.
Scientific Validity: Emphasizing content and construct validation in Phase IV.
Sustainable Impact: Promoting open practices and long-term governance in Phase V.
@@ 5.3 Future Directions: The Next Frontier of Medical Benchmark Research
Our work illuminates a clear path forward. Future efforts must move beyond the static, multiple-choice paradigm that currently dominates and pursue three critical directions:
Embracing Dynamic and Interactive Benchmarks. Future benchmarks must reflect the dynamic nature of clinical encounters by assessing sequential decision-making and information gathering. Pioneering works like MediQ (Li et al., 2024), which evaluates question-asking, and AgentClinic (Schmidgall et al., 2024), which uses multi-agent simulation, exemplify this necessary evolution toward evaluating core clinical reasoning skills (Trimble and Hamilton, 2016).
@@
Building a Collaborative Evaluation Ecosystem. To foster quality and transparency, the community should develop a living repository, akin to platforms like betterbench.stanford.edu (Reuel et al., 2024). Such a platform would enable continuous evaluation of benchmarks against a standardized framework like MedCheck. This would create a positive feedback loop, incentivizing higher-quality benchmarks and empowering all stakeholders to make more informed decisions about developing and deploying medical LLMs.
@@ 6 Conclusion
To address the lack of reliability in medical benchmarks, we introduced MedCheck, a lifecycle-oriented framework with 46 criteria. Our evaluation of 53 benchmarks revealed systemic deficiencies, including a "Clinical Disconnect" from real-world practice, a "Crisis of Data Integrity" due to contamination risks, and a "Systematic Neglect of Safety-Critical Capabilities." These issues underscore the need to shift from ad-hoc dataset creation to a more disciplined approach. MedCheck provides the essential toolkit for this shift, guiding the development of a rigorous infrastructure to move beyond the leaderboard and rethink medical LLM benchmarks, enabling the creation of genuinely safe and effective clinical AI.
@@
We acknowledge this study’s limitations. First, our analysis of 53 benchmarks, while extensive, is not exhaustive given the field’s rapid growth. Second, the scoring process, despite a rigorous protocol, retains a degree of subjectivity inherent in qualitative assessment. Third, our findings are based exclusively on public artifacts, which may not capture all unpublished development practices. Finally, the MedCheck framework itself is a snapshot of current best practices and will require future revisions to keep pace with evolving AI capabilities and ethical standards.
Acknowledgments
@@
In this appendix, we provide additional information about MedCheck. Appendix A lists the 53 evaluated medical LLM benchmarks. Appendix B provides extra quantitative analysis of the evaluation results. Appendix C shows the complete list of evaluation criteria.
Appendix A List of Evaluated Benchmarks
@@
We assessed the following 53 medical LLM benchmarks (alphabetical order):
  • AfriMed-QA Olatunji et al. (2025)
  • AgentClinic Schmidgall et al. (2024)
  • Asclepius Liu et al. (2025b)
  • BRIDGE Wu et al. (2025b)
  • CHBench Guo et al. (2024)
  • CheXpert Irvin et al. (2019)
  • CLIMB Dai et al. (2025)
  • CliMedBench Ouyang et al. (2024)
  • ClinicBench Liu et al. (2024a)
  • CMB Wang et al. (2024a)
  • cMedQA2 Zhang et al. (2018)
  • CMExam Liu et al. (2023)
  • COGNET-MD Panagoulias et al. (2024)
  • DataDEL Yang et al. (2024)
  • EHRNoteQA Kweon et al. (2024)
  • EndoBench Liu et al. (2025c)
  • ER-REASON Mehandru et al. (2025)
  • GMAI-MMBench Chen et al. (2024)
  • HeadQA Vilares and Gómez-Rodríguez (2019)
  • HealthBench Arora et al. (2025)
  • LLMEval-Med Zhang et al. (2025)
  • MedAgentBench Jiang et al. (2025)
  • MedAgentsBench Tang et al. (2025)
  • MedCalc Khandekar et al. (2024)
  • MedChain Liu et al. (2024b)
  • MedConceptsQA Shoham and Rappoport (2024)
  • MEDEC Abacha et al. (2025)
  • MedExQA Kim et al. (2024)
  • MEDIC Kanithi et al. (2024)
  • MediQ Li et al. (2024)
  • MedJourney Wu et al. (2025c)
  • MedMCQA Pal et al. (2022)
  • MedOdyssey Fan et al. (2025a)
  • MedQA Jin et al. (2021)
  • MedR-Bench Qiu et al. (2025)
  • MedRisk Liu et al. (2025a)
  • MedSafetyBench Han et al. (2024)
  • MedS-Bench Wu et al. (2025a)
  • MLEC-QA Li et al. (2021)
  • MMMU (Health & Medicine) Yue et al. (2024)
  • MVME (AI hospital) Fan et al. (2025b)
  • OmniMedVQA Hu et al. (2024)
  • PathMMU Sun et al. (2025b)
  • PathVQA He et al. (2020)
  • PMC-VQA Zhang et al. (2024)
  • PubMedQA Jin et al. (2019)
  • ReasonMed Sun et al. (2025a)
  • SLAKE Liu et al. (2021)
  • TrialPanorama Wang et al. (2025)
  • VQA-Med Ben Abacha et al. (2021)
  • VQA-RAD Lau et al. (2018)
  • webMedQA He et al. (2019)
  • XMedBench Wang et al. (2024b)
Appendix B Quantitative Results

### v2

@@ 4.4 Guideline for Benchmark Validity & Performance Verification
Findings: The Risk of Misdirected Progress. Formal scientific validation of benchmarks themselves is rare. While most benchmarks can differentiate models, only 54% (30 of 56) provide a compelling analysis of their content validity, and just 38% (21 of 56) are grounded in scenarios with high real-world authenticity. Without proper validation, a benchmark’s results are difficult to interpret. Poor content validity creates blind spots, while poor construct validity means the benchmark may not measure the claimed capability at all, leading to misleading conclusions and misdirected research efforts (Alaa et al., 2025). This lack of self-reflection risks optimizing for metrics that lack genuine clinical utility. To bridge this gap, developers must integrate clinician-in-the-loop validation, ensuring that automated evaluation metrics strictly align with physician preferences and patient safety outcomes.
@@ 4.5 Guideline for Documentation, Openness, & Governance
Findings: A Fragmented and Unsustainable Ecosystem. While most benchmarks provide public access to their assets (55 of 56), the governance required for sustainable impact is critically lacking. A significant 39% (22 of 56) do not specify a usage license, creating barriers to adoption. Most critically, 80% (45 of 56) have no clear maintenance plan, and 63% (35 of 56) lack a public feedback channel. This renders them "fire-and-forget" artifacts destined for obsolescence. This practice fosters a fragmented and unsustainable ecosystem, where effort is wasted on disposable artifacts rather than on building a lasting, reliable infrastructure for scientific progress (Arvan et al., 2022). To ensure longevity, the community must move toward institutional stewardship models, where host organizations provide explicit, long-term funding and technical commitments to prevent valuable benchmarks from becoming static, unmaintained relics.
@@ 5 Discussion
Our analysis of 56 medical benchmarks reveals a field at a critical crossroads. While research is active, systemic deficiencies threaten the validity and utility of this work.
@@ 5.1 Implications: The Urgent Need for a Paradigm Shift
The status quo is scientifically unsound and clinically irresponsible. The current trajectory fosters an "illusion of progress," where scores on clinically irrelevant or contaminated tasks mask a lack of genuine advancement. By failing to account for evaluation noise, the current paradigm risks optimizing models for pattern memorization rather than the deep, non-linear reasoning required for complex patient care (Laskar et al., 2024; Chang et al., 2023). This misdirects research efforts, misinforms stakeholders, and delays the responsible integration of AI into healthcare by creating brittle, biased systems that could harm patients (Obermeyer et al., 2019).
@@ 5.2 Our Guideline as a Catalyst for Change
Our five-phase lifecycle and 46-criteria checklist offer a practical toolkit and actionable roadmap for this shift, helping developers create higher-quality benchmarks and users assess existing ones. Adopting this framework can directly mitigate the critical issues identified in our study by promoting: Medical Grounding (mandating expert involvement and alignment with medical standards), Data Integrity (enforcing transparent sourcing, quality control, and contamination checks), Safety-Oriented Evaluation (requiring assessment of reasoning, robustness, and uncertainty), Scientific Validity (emphasizing content and construct validation), and Sustainable Impact (promoting open practices and long-term governance). Furthermore, it is essential to recognize that the multi-dimensional objectives within the MedCheck lifecycle are mutually reinforcing rather than fundamentally incompatible. For instance, stringent data privacy (Phase II) and open-source transparency (Phase V) act as co-existing requirements: rigorous de-identification is a strict prerequisite before a clinical dataset can be ethically open-sourced. High standards in one dimension fundamentally enable and secure practices in the others, creating a cohesive ecosystem for trustworthy AI evaluation.
@@ 5.3 Future Directions: The Next Frontier of Medical Benchmark Research
Future efforts should move beyond the static, multiple-choice paradigm and pursue three critical directions:
Embracing Dynamic and Interactive Benchmarks. Future benchmarks must reflect the dynamic nature of clinical encounters by assessing sequential decision-making and information gathering. Pioneering works like MediQ (Li et al., 2024) and AgentClinic (Schmidgall et al., 2024) exemplify this necessary evolution toward evaluating core clinical reasoning. Such interactive setups force models to actively elicit missing patient information, better mirroring the iterative and often ambiguous nature of real-world differential diagnosis (Trimble and Hamilton, 2016).
@@
Building a Collaborative Evaluation Ecosystem. To foster quality and transparency, the community should develop a living repository, akin to platforms like betterbench.stanford.edu (Reuel et al., 2024). Such a platform would enable continuous evaluation of benchmarks against a standardized framework like MedCheck. This would create a feedback loop incentivizing higher-quality benchmarks and more informed decisions.
@@ 6 Conclusion
To address the persistent reliability concerns in medical AI evaluation, we introduced MedCheck, a comprehensive, lifecycle-oriented assessment framework comprising 46 criteria. Our empirical evaluation of 56 benchmarks exposed systemic deficiencies across the field, namely a profound disconnect from real-world clinical practice, severe data integrity vulnerabilities stemming from unmitigated contamination risks, and a widespread neglect of safety-critical capabilities. These findings underscore the urgent need to transition from ad-hoc dataset curation toward a disciplined, engineering-focused approach. MedCheck serves as the foundational toolkit for this paradigm shift, guiding the community to move beyond superficial leaderboard rankings and rethink medical LLM benchmarks entirely, ultimately fostering the development of genuinely safe, trustworthy, and clinically effective artificial intelligence.
@@
We acknowledge this study’s limitations. First, our analysis of 56 benchmarks, while extensive, is not exhaustive given the field’s rapid growth. Second, the scoring process, despite a rigorous protocol, retains a degree of subjectivity inherent in qualitative assessment. Third, our findings are based exclusively on public artifacts, which may not capture all unpublished development practices. Finally, the MedCheck framework itself is a snapshot of current best practices and will require future revisions to keep pace with evolving AI capabilities and ethical standards. Specifically, as multimodal and autonomous medical agents emerge, future iterations must expand to rigorously audit multimodal clinical grounding and agentic safety.
@@
In this appendix, we provide additional information about MedCheck. Appendix A discusses clinically-grounded evaluation and reporting frameworks. Appendix B shows the benchmark searching and selecting procedure, and Appendix C details our evaluation setup. Appendix D outlines the classification of clinical and medical benchmarks. Appendix E lists the 56 evaluated medical LLM benchmarks, and Appendix F provides extra quantitative analysis of the evaluation results. Appendix H shows the complete list of evaluation criteria. Finally, Appendix I provides detailed scoring and explanations for a representative benchmark, and Appendix J presents an actionable diagnostic report example based on a clinical-oriented benchmark.
Appendix A Clinically-Grounded Evaluation and Reporting Frameworks
Beyond the benchmark datasets themselves, clinical informatics researchers have developed a parallel stream of work focused on establishing rigorous methodologies for AI model evaluation, validation, and transparent reporting. This research, published in top-tier medical and clinical informatics journals, emphasizes clinical utility, patient safety, and real-world applicability over leaderboard performance alone. A cornerstone of this ecosystem is the decades-long MIT and Harvard Medical School effort to develop, de-identify, and ethically share large-scale EHR datasets like MIMIC. The MIMIC project represents more than a data source—it embodies the principles of data curation, long-term maintenance, and community governance that are central to MedCheck’s framework, particularly in Phase II and Phase V.
Recent work emerging from this clinically-grounded tradition has produced several critical resources. The TRIPOD-LLM statement, for example, provides the first reporting guideline specifically for studies using LLMs, ensuring that methods and results are communicated transparently (Gallifant et al., 2025). While TRIPOD-LLM guides the reporting of a study, our MedCheck framework is a complementary tool designed to assess the quality of the benchmark used within that study. In parallel, comprehensive evaluation platforms like Med-HELM have been developed to assess LLMs on a wide array of clinically-grounded tasks using real EHR data. Med-HELM’s focus is on holistically evaluating the model’s capabilities, whereas MedCheck’s focus is on the prior, meta-level task of evaluating the measurement instrument itself.
The very premise of evaluating LLMs on medical tasks is predicated on the knowledge they contain. The landmark work by Singhal et al. (2023) was among the first to systematically demonstrate that LLMs do, in fact, encode a substantial amount of clinical knowledge. This finding underscores the urgency and importance of developing robust, clinically-valid benchmarks to accurately measure and safely harness this encoded knowledge, moving beyond exam-style questions to tasks that reflect the complexity of real-world clinical practice.
Appendix B Benchmark searching and selection
To ensure the objectivity and comprehensiveness of our evaluation, we implemented a systematic literature search and selection protocol divided into three stages: identification, screening, and eligibility. We first queried major academic databases—including Google Scholar, ACL Anthology, arXiv, and PubMed—for publications from January 2018 to July 2025. The search strategy employed Boolean combinations of keywords across subject, domain, and artifact type, specifically incorporating terms such as "medical LLM," "medical benchmark," "healthcare," "clinical," "clinical decision," and data-specific identifiers like "EHR" (electronic health record) and "EMR." This initial retrieval, focused on titles and abstracts to ensure relevance, yielded a total of 482 candidate publications after removing duplicates.
Subsequently, we applied stringent exclusion criteria to filter these candidates. We excluded studies strictly limited to traditional discriminative paradigms (e.g., BERT-based classification) that have not been adapted for evaluating the generative reasoning capabilities of LLMs. Furthermore, we removed benchmarks lacking publicly accessible data or code repositories essential for our reproducibility analysis, as well as studies that merely re-evaluated existing datasets without introducing novel contributions. This screening process narrowed the candidate pool to 85 potential benchmarks.
From this screened corpus, we selected the final set of 56 benchmarks based on specific criteria regarding impact and community adoption. Addressing the ambiguity of "widely used," we defined significance through quantitative metrics tailored to the publication timeframe. For established benchmarks published prior to 2023, we required a minimum citation count of 50 to verify community adoption. Conversely, for emerging benchmarks published in 2024 and 2025, we selected those that were either accepted by top-tier venues (e.g., ACL, NeurIPS, Nature Medicine) or represented a unique task category not covered by older benchmarks. This multi-stage protocol ensures that our analysis captures the most significant and representative contributions to the medical LLM landscape.
Appendix C Evaluation Setup
In the evaluation process of this study, the primary focus was on manual scoring conducted by domain experts, with the use of LLMs providing supplementary assistance. While LLM were strictly instructed to follow the evaluation criteria, they inevitably exhibited hallucinations. Thus, the manual evaluation process remained the cornerstone of the benchmark assessment, with LLM serving as a tool to facilitate expert review by quickly retrieving relevant information and providing references for scoring.
This section will first describe the manual evaluation process, outlining how human annotators were trained, calibrated, and involved in scoring. It will then explain how the LLM-as-judge approach was implemented to enhance efficiency and assist in the evaluation process.
C.1 Guideline for annotator
To ensure reliable and consistent scoring of benchmarks according to the proposed evaluation framework, we established a rigorous annotation protocol executed by three expert annotators. The process comprised two phases: (1) a preparation and calibration phase, and (2) a formal scoring phase, as detailed below.
Table 6: Overall Scoring Criteria
Score Description
0 The criterion is completely absent, severely lacking, or contains fundamental flaws.
Judgment basis: No relevant information; clearly incorrect information.
1 Initial attempt made, but incomplete, lacking details, methodologically crude, or unvalidated.
Judgment basis: Mentioned but not implemented; implemented but without evaluation; described ambiguously, making validation impossible.
2 Clear, complete, transparent, and aligned with current domain consensus or literature-recommended best practices.
Judgment basis: Explicit description + implementation + validation; citation of authoritative methods; open-source / reproducible / peer-reviewed / recognized by the community.
Preparation and Calibration Phase
One week prior to scoring, all three annotators were provided with the full evaluation criteria (Appendix  H). After confirming that each annotator had thoroughly reviewed and understood the criteria, we convened an online calibration meeting to accomplish the following objectives: (a) Jointly review the evaluation criteria; (b) Discuss and resolve discrepancies, with particular attention to clarifying ambiguous standards or dimensions with similar names, and establish consensus on terminology and definitions; (c) Harmonize overall scoring expectations (As shown in Table 6); (d) Clarify acceptable sources of evidence: scores must be grounded in publicly accessible documentation (e.g., main paper text, appendices, GitHub repositories, technical reports, official websites). If required information is absent, the item receives a default score of 0, with a comment stating “Information not provided.”
After each annotator completed scoring at least three benchmarks, a second alignment meeting was held to address emerging ambiguities or inconsistencies encountered during initial scoring. Based on this discussion, all previously submitted scores were reviewed and revised as necessary to maintain inter-annotator consistency.
To ensure reliability and directly address the limitation of single-annotator-per-case for the main corpus, we established a shared evaluation subset of 5 randomly selected benchmarks before the formal scoring phase. All three annotators independently scored this subset. We calculated the inter-annotator agreement using Fleiss’ Kappa, achieving a strong agreement score ($\kappa=0.78$), which indicates substantial consensus on the interpretation of the criteria.
Formal Scoring Phase
A total of 56 benchmarks were uniformly and randomly distributed among the annotators. Each annotator received a randomly ordered list of benchmarks, accompanied by the LLM-generated scores and corresponding explanations for each benchmark, and was instructed to manually evaluate 1–2 benchmarks per day, adhering to the following requirements:
(a) For each scored item, provide specific evidence in the comment field (As illustrated in the “Explanation” section of Appendix I); (b) Ignore authorship and institutional affiliations; avoid basing judgments on general impressions or prior familiarity with a benchmark; (c) Apply criteria strictly as defined—do not broaden or narrow the scope based on personal interpretation; (d) The LLM-generated scores and explanations are intended solely to serve as auxiliary aids—facilitating rapid information retrieval and offering preliminary scoring suggestions—and must not constitute the primary basis for evaluation. Their content must be rigorously verified for factual accuracy before being considered as reference material; (e) For items with insufficient or ambiguous information, assign a score of 0, 1, or 2 as appropriate, but include a brief justification in the comments (e.g., “The paper mentions data cleaning but does not detail the procedure”) and prefix the note with “[Uncertain]”; (e) For borderline cases where a benchmark’s documentation is vague or implicitly implies a feature, annotators were strictly instructed to default to the lower score to penalize the lack of transparency. A higher score was only awarded if concrete, verifiable evidence (e.g., explicit code snippets, specific GitHub commits, or dedicated appendix sections) could be explicitly cited in the comments. Any remaining uncertainties (flagged with "[Uncertain]") were brought to the weekly expert panel meetings, where consensus was reached through rigorous group discussion and majority voting; (f) Daily monitoring of individual annotator score distributions was conducted to detect and mitigate systematic bias (e.g., consistently high or low scoring tendencies).
Throughout the process, any issues or discrepancies in scoring were resolved through weekly consensus meetings among the expert panel, culminating in a final agreed-upon score for each criterion. This rigorous protocol substantially mitigated the well-documented susceptibility of manual scoring to various confounding factors—such as subjective bias, cognitive load, and order effects—and ensured that the evaluation process was maximally transparent, consistent, and reproducible.
C.2 LLM-as-Judge Scoring
While the primary evaluation in MedCheck is conducted manually by domain experts, we explored the use of LLM-as-judge scoring during the preliminary scoring phase to balance scalability with accuracy.
We include our evaluation criteria, the paper text, and documentation of medical benchmarks in the prompt. We then use GPT-4o-Search-Preview with temperature = 0 to generate scores in JSON format. To ensure the LLM produces reasonable scores based on our criteria, we explicitly instructed it to strictly adhere to the evaluation guidelines and provide an explanation for each score. Additionally, to guarantee the LLM has access to up-to-date public documentation, we enabled web search functionality, allowing it to retrieve supplementary information (e.g., the benchmark’s official website, GitHub repository, etc.) when necessary. During the manual scoring process, the panel of domain experts refers to the LLM-generated scores and explanations, but they must verify the accuracy of the LLM-provided explanations and ultimately assign the final scores.
To address potential concerns regarding confirmation bias from LLM-generated scores, we explicitly frame our approach as an "Assisted-Verification" design. In the context of auditing extensive and complex benchmark documentation, independent "blind" human annotation often suffers from high false-negative rates due to cognitive fatigue and information overload (i.e., annotators inadvertently missing existing evidence). In our protocol, the LLM functions strictly as an information retrieval assistant to surface relevant sections and ensure high recall. The human expert then rigorously verifies the factual existence of the cited documentation to ensure high precision. Because the assessment relies on objective verification of factual evidence rather than subjective judgment, the risk of bias is minimized while the accuracy of evidence retrieval is significantly improved.
In summary, the evaluation process relied fundamentally on the expertise of human annotators to ensure the accuracy and consistency of the final scores. LLMs played a strictly auxiliary role—expediting data retrieval and providing preliminary scoring suggestions—while the ultimate judgment remained firmly in the hands of domain experts. This hybrid methodology synergistically combines the rigor and reliability of manual evaluation with the efficiency and scalability of LLMs, thereby ensuring an evaluation framework that is both comprehensive and practically viable.
Appendix D Classification of Clinical and Medical Benchmarks
We classify the 56 evaluated benchmarks into two distinct categories: Clinical (42 benchmarks, 75%) and Medical (14 benchmarks, 25%). These categories exhibit a fundamental distinction in their objectives, application scenarios, and data sources. We argue that evaluating a benchmark from one category using the criteria intended for the other would be methodologically flawed. Therefore, to ensure a fair and context-aware assessment, the criteria used to define each category are detailed below.
  • Clinical Benchmarks (42/56, 75%): These benchmarks primarily evaluate the application of LLMs in real-world clinical workflows.
    • Objective: Assess capabilities like processing EHRs, clinical reasoning with dynamic or incomplete information, and supporting diagnostic decisions.
    • Scenario: Simulate authentic clinical encounters, such as patient consultations, risk prediction, or treatment planning.
    • Data Source: Primarily use real-world data, including EHRs, clinical case notes, and doctor-patient dialogues.
  • Medical Benchmarks (14/56, 25%): These benchmarks focus on assessing an LLM’s foundational medical knowledge and theoretical reasoning.
    • Objective: Test the model’s mastery of established medical facts and concepts.
    • Scenario: Typically involve standardized, knowledge-based tasks, often in a multiple-choice question format.
    • Data Source: Primarily use academic materials, such as medical exam questions, textbooks, and research literature.
Appendix E List of Evaluated Benchmarks
@@
We assessed the following 56 benchmarks, categorized as either Clinical (C) or Medical (M) (alphabetical order):
  • AfriMed-QA Olatunji et al. (2025) (C)
  • AgentClinic Schmidgall et al. (2024) (C)
  • Asclepius Liu et al. (2025b) (C)
  • BRIDGE Wu et al. (2025b) (C)
  • CHBench Guo et al. (2024) (C)
  • CheXpert Irvin et al. (2019) (C)
  • CLIMB Dai et al. (2025) (C)
  • CliMedBench Ouyang et al. (2024) (C)
  • ClinicBench Liu et al. (2024a) (C)
  • CMB Wang et al. (2024a) (M)
  • cMedQA2 Zhang et al. (2018) (C)
  • CMExam Liu et al. (2023) (M)
  • COGNET-MD Panagoulias et al. (2024) (M)
  • DataDEL Yang et al. (2024) (C)
  • DR.BENCH Gao et al. (2023) (C)
  • EHRNoteQA Kweon et al. (2024) (C)
  • EndoBench Liu et al. (2025c) (C)
  • ER-REASON Mehandru et al. (2025) (C)
  • GMAI-MMBench Chen et al. (2024) (C)
  • HeadQA Vilares and Gómez-Rodríguez (2019) (M)
  • HealthBench Arora et al. (2025) (C)
  • LCD Yoon et al. (2025) (C)
  • LLMEval-Med Zhang et al. (2025) (C)
  • MedAgentBench Jiang et al. (2025) (C)
  • MedAgentsBench Tang et al. (2025) (M)
  • MedCalc Khandekar et al. (2024) (C)
  • MedChain Liu et al. (2024b) (C)
  • MedConceptsQA Shoham and Rappoport (2024) (M)
  • MEDEC Abacha et al. (2025) (C)
  • MedExQA Kim et al. (2024) (M)
  • MEDIC Kanithi et al. (2024) (C)
  • MediQ Li et al. (2024) (C)
  • MedJourney Wu et al. (2025c) (C)
  • MedMCQA Pal et al. (2022) (M)
  • MedOdyssey Fan et al. (2025a) (C)
  • MedQA Jin et al. (2021) (M)
  • MedR-Bench Qiu et al. (2025) (C)
  • MedRisk Liu et al. (2025a) (C)
  • MedSafetyBench Han et al. (2024) (C)
  • MedS-Bench Wu et al. (2025a) (C)
  • MIMIC-IV-ED Xie et al. (2022) (C)
  • MLEC-QA Li et al. (2021) (M)
  • MMMU (Health & Medicine) Yue et al. (2024) (C)
  • MVME (AI hospital) Fan et al. (2025b) (C)
  • OmniMedVQA Hu et al. (2024) (C)
  • PathMMU Sun et al. (2025b) (C)
  • PathVQA He et al. (2020) (M)
  • PMC-VQA Zhang et al. (2024) (C)
  • PubMedQA Jin et al. (2019) (M)
  • ReasonMed Sun et al. (2025a) (M)
  • SLAKE Liu et al. (2021) (C)
  • TrialPanorama Wang et al. (2025) (C)
  • VQA-Med Ben Abacha et al. (2021) (C)
  • VQA-RAD Lau et al. (2018) (C)
  • webMedQA He et al. (2019) (C)
  • XMedBench Wang et al. (2024b) (M)
Appendix F Quantitative Results